OpenStudy (crashonce):

Prove cos(A+B)cos(A-B) = cos^2(A) - sin^2(B)

OpenStudy (crashonce):

@zepdrix @ParthKohli @No.name

OpenStudy (crashonce):

@ikram002p

OpenStudy (ikram002p):

ok , so expand !

OpenStudy (ikram002p):

cos (A+B)= ? cos(A-b) = ?

OpenStudy (crashonce):

RHS? which way tho? i could expand two angles to singles or elimnate the product

OpenStudy (crashonce):

so as in (cosAcosB-sinAsinB)(cosAcosB+sinAsinB)?

OpenStudy (crashonce):

@ikram002p

OpenStudy (ikram002p):

\(\cos (A+B )=\cos A \cos B -\sin A \sin B\) \(\cos (A-B )=\cos A \cos B +\sin A \sin B\) thus \((\cos (A+B ))(\cos (A-B ) )=(\cos A \cos B -\sin A \sin B)(\cos A \cos B +\sin A \sin B)\)

OpenStudy (ikram002p):

oh yeah @CrashOnce what u did is correct :) now FOIL expand

OpenStudy (crashonce):

(cosAcosB)^2-(sinAsinB)^2

OpenStudy (ikram002p):

correct ! since (a+b)(a-b)=a^2-b^2

OpenStudy (crashonce):

but that isnt cos^2A-sin^2B

OpenStudy (ikram002p):

hehe , looking for other identity :P

Parth (parthkohli):

Alternatively, use the identity\[2\cos A \cos B = \cos(A - B) - \cos(A + B)\]In this case,\[\cos(A + B)\cos(A -B ) = \frac{1}{2}\left(\cos (2B)-\cos(2A)\right)\]\[= \frac{1}{2}\left(1 - 2\sin^2 B + \cos^2 A - 1 \right)\]\[= \cdots\]

Parth (parthkohli):

Correction:\[= \frac{1}{2}\left(1 - 2\sin^2 B + 2\cos^2 A - 1 \right) \]

OpenStudy (ikram002p):

nice !